# Robust Multi-Task Learning with Excess Risks

### v2

@@ C.1 Implementation
We implement our algorithm using hard parameter sharing, where all tasks share a feature extractor and have task-specific heads. For feature extractors, we use a two-layer CNN for MultiMNIST, ResNet-18 [He et al., 2016] for Office-Home, and SegNet [Badrinarayanan et al., 2017] for NYUv2. For task-specific heads, we use two-layer CNNs for NYUv2, and MLP for all other datasets. We standardize all datasets to ensure zero mean and unit variance, as excess risks are sensitive to the scale of tasks. The details are as follows.
@@
On NYUv2, we follow the implementation of Liu et al. [2019]. Since the dataset have limited data, making it prone to overfitting, we use data augmentation as suggested by Liu et al. [2019]. For the feature extractor, we use the SegNet architecture with four convolutional encoder layers and four convolutional decoder layers. For each of the three tasks, we use two convolutional layers as the task-specific heads. We use a weight decay of 1e-3 and the learning rate is 1e-4.
@@
For the implementation of baselines, we use the code from Lin and Zhang [2023] and Navon et al. [2022].
@@ C.2 More Results
SARCOS [Vijayakumar and Schaal, 2000] presents an inverse dynamics problem for a robot arm with seven degrees of freedom. The task is to perform multi-target regression that uses 21 attributes (7 joint positions, 7 joint velocities, 7 joint accelerations) to predict the corresponding 7 joint torques. The noise is injected into the first two joint torques.

### v3

@@ C.1 Implementation
We implement our algorithm using hard parameter sharing, where all tasks share a feature extractor and have task-specific heads. For feature extractors, we use a two-layer CNN for MultiMNIST, ResNet-18 (He et al., 2016) for Office-Home, and SegNet (Badrinarayanan et al., 2017) for NYUv2. For task-specific heads, we use two-layer CNNs for NYUv2, and MLP for all other datasets. We standardize all datasets to ensure zero mean and unit variance, as excess risks are sensitive to the scale of tasks. The details are as follows.
@@
On NYUv2, we follow the implementation of Liu et al. (2019). Since the dataset have limited data, making it prone to overfitting, we use data augmentation as suggested by Liu et al. (2019). For the feature extractor, we use the SegNet architecture with four convolutional encoder layers and four convolutional decoder layers. For each of the three tasks, we use two convolutional layers as the task-specific heads. We use a weight decay of 1e-3 and the learning rate is 1e-4.
@@
For the implementation of baselines, we use the code from Lin & Zhang (2023) and Navon et al. (2022).
@@ C.2 More Results
SARCOS (Vijayakumar & Schaal, 2000) presents an inverse dynamics problem for a robot arm with seven degrees of freedom. The task is to perform multi-target regression that uses 21 attributes (7 joint positions, 7 joint velocities, 7 joint accelerations) to predict the corresponding 7 joint torques. The noise is injected into the first two joint torques.